OpenStudy (anonymous):

a geometric progression for which the common ratio is positive has a second term 18 and a fourth term of 8 find 1. the first term and the common ratio progression 2. the sum to infinity of the progression show how you did

OpenStudy (tkhunny):

a = first term r = common ratio 1st term = a 2nd term = ar = 18 3rd term = ar^2 4th term = ar^3 = 8 ... Sum = a/(1-r) Go!!

OpenStudy (anonymous):

substitute in sn=a(1-r)^n/1-r ?

OpenStudy (tkhunny):

No, that is not quite the right formula for a FINITE geometric series. I provided the right one for an INFINITE geometric series. Your task is to find 'a' and 'r'.

OpenStudy (anonymous):

the formula is for geometric progression not sum also i dont any other ways of findin a and r

OpenStudy (tkhunny):

You must look for clues rather than throwing your hands in the air. The problem statement provides these clues. ar = 18 ar^3 = 8

OpenStudy (anonymous):

i throw my hand in the air saying ayoo gotta let go also i was saying a= 18/r and then substitute this in formula in place of a for finding ratio

OpenStudy (tkhunny):

Very good. What do you get for 'a' and 'r'?

OpenStudy (anonymous):

solving in progress

OpenStudy (anonymous):

so far i am @ 18r^2 - 18r = 18^2 - 18r^3 can u tell if i am doin it right ?

OpenStudy (tkhunny):

What happened to the lovely a = 18/r from the 2nd term?? Substituting this value into the 4th term: (18/r)r^3 = 8 18r^2 = 8 r^2 = 8/18 = 4/9 Now what?

OpenStudy (anonymous):

damn i subbed in formula for progression ill take square root and go over you method again

OpenStudy (tkhunny):

Careful with "take the square root". There are TWO solutions to r^2 = 4/9.

OpenStudy (anonymous):

its + and - same answer but question says its positive so

OpenStudy (tkhunny):

Bonus points for referring to the problem statement to resolve that conflict! Excellent work.

OpenStudy (tkhunny):

So, r = ?????

OpenStudy (anonymous):

i got 2/3 but i think its wrong

OpenStudy (tkhunny):

No, r = 2/3. This, however is not the end. a = 18/r = ?????

OpenStudy (anonymous):

27

OpenStudy (anonymous):

is it right ?

OpenStudy (tkhunny):

Okay, we are ready for the sum to infinity!

OpenStudy (tkhunny):

\(\dfrac{a}{1-r} = \dfrac{27}{1 - (2/3)} =\;??\)